I'm down to my last 3 problems, but I don't know how to solve them.
1. 12x - 1 - 11x - 1 = -6

2. 12 - 16 = 4y + 6 - 3y

3. 6x - 6 - 5x + 1 = -2 + 4

Answer by Abbey(339)

1. 12x - 1 - 11x - 1 = -6
combine like terms
x-2=-6
add 2 to both sides
x=-4
2. 12 - 16 = 4y + 6 - 3y
combine like terms
4=y+6
subtract 6 from both sides:
-2=y
3. 6x - 6 - 5x + 1 = -2 + 4
combine like terms
x-5=2
add 5 to both sides:
x=7